OpenStudy (anonymous):

3/8 divided by 1/2= i got 3/4 but it maybe 7/12???

OpenStudy (zzr0ck3r):

\[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

i think 100% its 3/4

OpenStudy (zzr0ck3r):

so \[\frac{\frac{3}{8}}{\frac{1}{2}}=\frac{3}{8}*\frac{2}{1}=\frac{6}{8}=\frac{3}{4}\]you are correct

OpenStudy (anonymous):

ok thanks